# Privacy Risks and Preservation Methods in Explainable Artificial Intelligence: A Scoping Review

### v2

@@ 4.1 Threat model
To discuss the threat model of XAI, we consider a target AI model, f, with a corresponding explanation function, $\phi$. For an input query, x, the system generates an output, f(x), such as prediction, classification or an artifact of a Gen-AI system, and a corresponding explanation, $\phi$(x) (Figure 5). Though the explanation interface is made available for the end-users of the system, adversaries can also secure different levels of access to the XAI system during the stages of the AI lifecycle (Shahriar et al., 2023). In white-box access, adversaries possess information about the model internals such as architecture and hyperparameters (Liu et al., 2022a; Zhang et al., 2023). In black-box access, adversaries access the input/output of the system (Liu et al., 2022a) without any knowledge of the training process or model internals (Rigaki & Garcia, 2023). Any intermediate access between white-box and black-box is referred to as gray-box (Jegorova et al., 2022). The adversaries may act as passive observers (Jegorova et al., 2022) and use the model outputs for launching privacy attacks, or they may actively interfere in the training process of the model (Nasr et al., 2019).
@@ 4.2 Intentional privacy leakage
This subsection reviews intentional risks in the form of privacy attacks launched by adversaries. As XAI systems are fundamentally AI models augmented with explainability features, they remain susceptible to malicious threats that affect conventional AI models. Prior research has identified security attacks, such as evasion and poisoning (Pitropakis et al., 2019), that compromise the integrity of AI. Adversarial attacks on XAI such as input manipulations (Dombrowski et al., 2019; Zhang et al., 2020), model manipulations (Heo et al., 2019) and explanation-aware backdoors (Noppel et al., 2023) are also discussed in current literature. However, the present article focusses on privacy attacks that aim to compromise the personal data of individuals or the confidentiality of the underlying model. In the XAI context, model explanations further aid (Zhao et al., 2021) the identification or exposure of personal information of individuals or the intellectual property of the model owner. Table LABEL:tab:table2 provides an overview of these intentional privacy risks through XAI and the studies addressing them.
@@ 4.2.1 Membership inference
This is a privacy risk of identification of an individual in the training set of a model (Shokri et al., 2017; Zarifzadeh et al., 2024) (Figure 6). An adversary can execute this attack with black-box or white-box access to the model (Veale et al., 2018) after it has been deployed. A membership inference model can be expressed as the following binary classifier (Kuppa & Le-Khac, 2021) when the model output, $f(x)$, is available:
@@
The seminal work on membership inference in feature-based and example-based XAI systems was presented by Shokri et al. (2020; 2021). The study used various backpropagation and perturbation methods to show the vulnerability of feature-based systems. The proposed attack used variances in the prediction and explanation vectors to differentiate between members and non-members. Liu et al. (2024) introduced a membership inference on feature-based XAI using model performance and robustness metrics. The study observed higher loss in confidence on perturbation of important features for members and utilized this observation in training an attack model, in addition to using the performance loss from the model. Ma et al. (2024) extended membership inference to label-only settings on Shapley value explanations. This method, which builds on earlier work on label-only attacks using hard prediction labels (Choquette-Choo et al., 2021), improved neighbourhood sampling using explanations thus reducing the number of queries.
@@
In the example-based category, Shokri et al. (2020; 2021) investigated influence functions on logistic regression models. Since influence functions generate explanations in the form of actual datapoints, the study observed that attackers could obtain certainty about membership, thus leading to stronger attacks. More recently, Cohen & Giryes (2024) considered self-influence functions instead, that show the influence of a datapoint on its own prediction. The proposed attack required white-box access to the target model parameters, activations, and gradients. The selection of an appropriate threshold range for self-influence scores associated with members was critical for this attack and was achieved by maximizing the balanced accuracy on the training set.
@@
Kuppa & Le-Khac (2021) used a different type of example-based explanation, namely, counterfactuals, for membership inference. The authors trained shadow models using counterfactual samples and auxiliary datasets. A threshold on the difference in predictions of the attack and target models was used to determine membership. Pawelczyk et al. (2023) also targeted counterfactuals and proposed two types of attacks. The first relied on the distances between datapoints and their counterfactuals to differentiate between members and non-members. The second used a loss-based approach using a likelihood-ratio test (Carlini et al., 2022) that improved the attack.
@@
Interpretable models using decision trees, and surrogate models created using Trepan algorithm (Craven & Shavlik, 1995), were evaluated for membership inference by Naretto et al. (2022). The study also examined the effect of overfitting on the attack. The success of membership inference was determined to be higher on both interpretable and surrogate models compared to black-box models. Further, surrogates of overfitted models exhibited higher susceptibiliy to the attack than those derived from well-regularized models.
@@
Membership inference attacks in machine learning models have been explored extensively in existing literature (Hu et al., 2022) and attack strategies have exploited confidence scores and predictions (Shokri et al., 2017). However, the recent attacks that exploit explanations suggest that XAI interfaces provide a new avenue for adversaries to launch this attack. Such attacks have targeted feature-based, example-based, and interpretable (including surrogates) XAI methods. The effectiveness of the attack is influenced by factors such as dataset type (Shokri et al., 2021), dimension (Pawelczyk et al., 2023; Shokri et al., 2021), model architecture (Shokri et al., 2021) and overfitting (Pawelczyk et al., 2023). Some attacks have proven effective in the absence of knowledge of the training dataset or target architectures (Liu et al., 2024), underscoring their practical threat potential.
@@
While interpretable models are often recommended as surrogates for explaining black-box models (McDermid et al., 2021), as demonstrated by these attacks, the layer of interpretability can introduce a backdoor to the target system and lead to privacy leaks (Naretto et al., 2022). In the example-based category, influence functions expose data instances, particularly outliers, due to their distinct characteristics and higher influence on the training process (Shokri et al., 2021). Among feature-based methods, those using perturbations exhibit higher resilience to membership inference due to use of out-of-distribution points, however, this can also result in reduced explanation fidelity ((Shokri et al., 2021). Conversely, feature-based methods with better explanation quality are also found to be susceptible to higher leakage (Liu et al., 2024) suggesting a conflict between privacy and utility.
@@ 4.2.2 Model inversion
This category of privacy risk can result in reconstruction of datapoints partly or completely from outputs (Zhang et al., 2023) (Figure 7). These attacks can be conducted with black-box or white-box access to the model (Fredrikson et al., 2015; Veale et al., 2018) after it has been deployed in production systems. Attribute inference is a type of data reconstruction that can determine the values of certain attributes, generally those sensitive to individuals (Yeom et al., 2018) from outputs. When the model output, $f(x)$, is used for inversion, the inference model can be expressed as follows:
@@
Model inversion attacks have been documented in XAI on example-based, feature-based, and interpretable systems. Shokri et al. (2020; 2021) demonstrated a data reconstruction attack on influence functions in logistic regression models and found the attack dependent on data dimensionality. The authors designed different heuristics for low and high dimension data to improve coverage and efficiently recover more training points. Goethals et al. (2023) showed an explanation linkage attack using native counterfactuals generated from actual instances of the training data. The attack demonstrated the vulnerability of counterfactuals in leaking private attributes.
@@
Private images were found to be susceptible to recovery through saliency maps by Zhao et al. (2021) leading to inadvertent exposure. The study found XAI systems that provided class-specific multiple explanations particularly prone to leakage. The authors also used attention transfer to highlight similar risks for non-explanation models. Other studies (Duddu & Boutet, 2022; Luo et al., 2022) have focused on attribute inference of tabular data using feature-based XAI. The former trained attack models using predictions and explanations to infer sensitive features. The latter used Shapley values and effectively executed the attack with limited number of queries on cloud ML services. Toma & Kikuchi (2024) further showed that the efficacy of the proposed attack was dependent on the combination of black-box architecture and XAI method. Their findings indicate that linear models using Shapley values were particularly vulnerable to attribute inference.
@@
Ferry et al. (2024) designed a probabilistic white-box attack applicable to transparent models, such as decision trees and rule lists, and quantified the information about the training data contained in the model. The work found that models built using greedy algorithms leak more information compared to those built using optimal strategies. The authors also observed the attack’s capacity for misuse in launching other inference attacks such as membership and property inference.
@@
The above attacks from current literature demonstrate the potential of misuse of explanations APIs, resulting in a two-way flow (Veale et al., 2018) where inputs can be determined from outputs. Model explanations provide an effective attack surface compared to predictions (Duddu & Boutet, 2022; Zhao et al., 2021), indicating the contradiction between the need for explanations in Trustworthy AI and protecting privacy (Zhao et al., 2021). Data reconstruction attacks impact active users of AI systems, thus putting end-users at risk (Zhao et al., 2021) and having a higher impact. In certain proposed model inversion attacks, sensitive attributes can be retrieved from models trained on non-sensitive attributes (Duddu & Boutet, 2022) while other proposed attacks demonstrate higher leakage from more important attributes (Luo et al., 2022) and recovery of entire training datasets (Shokri et al., 2021). In addition, the above works highlight the misuse of XAI techniques even for models that do not provide explanations (Zhao et al., 2021).
@@
A tension is also found between preserving privacy and maintaining utility of the XAI system. For instance, synthetic counterfactuals created by perturbing actual samples are shown to provide resilience to inversion in comparison to native counterfactuals. However, their usage is found to affect the plausibility and runtime of explanations (Goethals et al., 2023), suggesting degrading utility. Similarly, the use of multiple diverse explanations are usually recommended (Vo et al., 2023) for improving understandibility of explanations, however they are also found to contribute to leakage of privacy. Consequently restricting the access to explanation APIs has been suggested as a countermeasure (Zhao et al., 2021), however such restrictions may reduce the utility to end-users.
@@ 4.2.3 Model extraction
In this attack, the target model is replicated to a significant degree of accuracy and fidelity (Jegorova et al., 2022), thus breaching the confidentiality and the intellectual property of the model owner (Figure 8). In a typical model extraction attack, the adversary has black-box access to a deployed victim model and uses an unlabeled dataset to query it, thus building an attack dataset (Yan et al., 2023b) for cloning the model. In contrast, data-free model extraction leverages generative models to synthesize the datasets, which is advantageous when the input data is difficult to obtain (Miura et al., 2024). When the model output, $f(x)$, is utilized for this attack, the extraction of the target model function, $f^{\prime}$, can be expressed as follows:
@@
Since the extracted models can further leak personal data through membership inference and model inversion (Song et al., 2017), this attack is usually used as a starting point for initiating other types of attacks (Aïvodji et al., 2020; Miura et al., 2024).
@@
Model extraction attacks using explanations have been proposed across feature-based and example-based XAI methods. In the seminal work on the topic, gradient explanations, in the form of saliency maps, were found to be vulnerable by Milli et al. (2019). The use of explanations improved the attack by reducing the number of queries compared to attacks relying solely on model predictions. Miura et al. (2024) also leveraged gradient-based explanations but used data-free approach to train generative models for creating the attack dataset. The inclusion of explanations improved the quality of the generative model, and the accuracy of the cloned model improved with the diversity of the generated samples. Similarly, Yan et al. (2023a) employed data-free extraction wherein explanation loss was used to guide the generative model. Accuracy of the cloned model was improved by matching the victim model’s predictions and explanations.
@@
A different approach of extraction on feature-based XAI, used multitask learning to learn both classification and explanation tasks of the victim model (Yan et al., 2022). Further, a model agnostic technique on gradient and pertubation-based XAI (Yan et al., 2023b), showed that explanations provided auxiliary information that enabled more efficient attacks, reducing the query budget compared to prediction-only strategies. The attack could also be applied to non-explanation models and achieved accuracy equivalent to those of explanation models.
@@
Besides the above extraction attacks targeting various feature-based XAI, from the example-based category, counterfactuals have been mainly targeted for this attack. In an extraction attack proposed by Aïvodji et al. (2020), they were used to approximate the decision boundary of the victim model with high accuracy and fidelity under low query budgets. Multiple and diverse counterfactuals were found to aid the extraction process by divulging additional information to adversaries. An improvisation of the attack, to reduce the number of queries further, mitigate decision boundary shift and achieve higher agreement with the victim model, was proposed by Wang et al. (2022). The method used the original counterfactual explanation with its own counterfactual as training pairs, to extract additional datapoints to train the cloned model. In another approach, Kuppa & Le-Khac (2021) proposed iterative querying of the victim model to capture the training data distribution. The method utilized distillation loss to transfer knowledge from the victim to the cloned model and was found to be successful due to the optimization of various properties such as diversity, proximity, feasibility, and sparsity.
@@
As demonstrated by the aforementioned attacks, explanation-based extraction attacks offer substantial advantages over traditional prediction-only approaches by facilitating model replication with reduced number of queries (Milli et al., 2019; Miura et al., 2024). The reduction in the number of queries benefits the adversary, especially in pay-by-query models. Certain attacks are also possible with partial knowledge of the data distribution (Aïvodji et al., 2020) or in absence of overlap between the attack and training datasets (Yan et al., 2022). In addition, in scenarios where attackers do not possess the input datasets, data-free extraction attacks are possible and the use of explanations is shown to improve the attack accuracy (Miura et al., 2024). Moreover, the diversity of the generated input datasets in such attacks is found to improve the accuracy of the cloned models (Miura et al., 2024).
@@
In addition to the direct threat to explanation models, XAI techniques can be misused for extraction of non-explanation models (Yan et al., 2023b). The use of diverse explanations, intended to build user trust in explanations, can lead to further leakage of privacy (Aïvodji et al., 2020). Similarly, the optimization of counterfactuals to satisfy various properties to improve explanation quality, can reveal information to adversaries about class-specific decision boundaries thus aiding the attack (Kuppa & Le-Khac, 2021) and leading to the conflict of explainability and privacy with utility.

### v3

@@ 4.1 Threat model
To discuss the threat model of XAI, we consider a target AI model, f, with a corresponding explanation function, $\phi$. For an input query, x, the system generates an output, f(x), such as prediction, classification or an artifact of a Gen-AI system, and a corresponding explanation, $\phi$(x) (Figure 5). Though the explanation interface is made available for the end-users of the system, adversaries can also secure different levels of access to the XAI system during the stages of the AI lifecycle (shahriarSurveyPrivacyRisks2023). In white-box access, adversaries possess information about the model internals such as architecture and hyperparameters (liuWhenMachineLearning2022; zhangSurveyPrivacyInference2023). In black-box access, adversaries access the input/output of the system (liuWhenMachineLearning2022) without any knowledge of the training process or model internals (rigakiSurveyPrivacyAttacks2023). Any intermediate access between white-box and black-box is referred to as gray-box (jegorovaSurveyLeakagePrivacy2022). The adversaries may act as passive observers (jegorovaSurveyLeakagePrivacy2022) and use the model outputs for launching privacy attacks, or they may actively interfere in the training process of the model (nasrComprehensivePrivacyAnalysis2019).
@@ 4.2 Intentional privacy leakage
This subsection reviews intentional risks in the form of privacy attacks launched by adversaries. As XAI systems are fundamentally AI models augmented with explainability features, they remain susceptible to malicious threats that affect conventional AI models. Prior research has identified security attacks, such as evasion and poisoning (pitropakisTaxonomySurveyAttacks2019), that compromise the integrity of AI. Adversarial attacks on XAI such as input manipulations (dombrowski2019explanations; zhang2020interpretable), model manipulations (heo2019fooling) and explanation-aware backdoors (noppel2023backdoor) are also discussed in current literature. However, the present article focusses on privacy attacks that aim to compromise the personal data of individuals or the confidentiality of the underlying model. In the XAI context, model explanations further aid (zhaoExploitingExplanationsModel2021) the identification or exposure of personal information of individuals or the intellectual property of the model owner. Table LABEL:tab:table2 provides an overview of these intentional privacy risks through XAI and the studies addressing them.
@@ 4.2.1 Membership inference
This is a privacy risk of identification of an individual in the training set of a model (shokriMembershipInferenceAttacks2017; zarifzadehLowCostHighPowerMembership2024) (Figure 6). An adversary can execute this attack with black-box or white-box access to the model (vealeAlgorithmsThatRemember2018) after it has been deployed. A membership inference model can be expressed as the following binary classifier (kuppaAdversarialXAIMethods2021) when the model output, $f(x)$, is available:
@@
The seminal work on membership inference in feature-based and example-based XAI systems was presented by shokriExploitingTransparencyMeasures2020; shokriPrivacyRisksModel2021. The study used various backpropagation and perturbation methods to show the vulnerability of feature-based systems. The proposed attack used variances in the prediction and explanation vectors to differentiate between members and non-members. liuPleaseTellMe2024 introduced a membership inference on feature-based XAI using model performance and robustness metrics. The study observed higher loss in confidence on perturbation of important features for members and utilized this observation in training an attack model, in addition to using the performance loss from the model. maLabelOnlyMembershipInference2024 extended membership inference to label-only settings on Shapley value explanations. This method, which builds on earlier work on label-only attacks using hard prediction labels (choquette-chooLabelOnlyMembershipInference2021), improved neighbourhood sampling using explanations thus reducing the number of queries.
@@
In the example-based category, shokriExploitingTransparencyMeasures2020; shokriPrivacyRisksModel2021 investigated influence functions on logistic regression models. Since influence functions generate explanations in the form of actual datapoints, the study observed that attackers could obtain certainty about membership, thus leading to stronger attacks. More recently, cohenMembershipInferenceAttack2024 considered self-influence functions instead, that show the influence of a datapoint on its own prediction. The proposed attack required white-box access to the target model parameters, activations, and gradients. The selection of an appropriate threshold range for self-influence scores associated with members was critical for this attack and was achieved by maximizing the balanced accuracy on the training set.
@@
kuppaAdversarialXAIMethods2021 used a different type of example-based explanation, namely, counterfactuals, for membership inference. The authors trained shadow models using counterfactual samples and auxiliary datasets. A threshold on the difference in predictions of the attack and target models was used to determine membership. pawelczykPrivacyRisksAlgorithmic2023 also targeted counterfactuals and proposed two types of attacks. The first relied on the distances between datapoints and their counterfactuals to differentiate between members and non-members. The second used a loss-based approach using a likelihood-ratio test (carliniMembershipInferenceAttacks2022) that improved the attack.
@@
Interpretable models using decision trees, and surrogate models created using Trepan algorithm (cravenExtractingTreeStructuredRepresentations1995), were evaluated for membership inference by narettoEvaluatingPrivacyExposure2022. The study also examined the effect of overfitting on the attack. The success of membership inference was determined to be higher on both interpretable and surrogate models compared to black-box models. Further, surrogates of overfitted models exhibited higher susceptibiliy to the attack than those derived from well-regularized models.
@@
Membership inference attacks in machine learning models have been explored extensively in existing literature (huMembershipInferenceAttacks2022) and attack strategies have exploited confidence scores and predictions (shokriMembershipInferenceAttacks2017). However, the recent attacks that exploit explanations suggest that XAI interfaces provide a new avenue for adversaries to launch this attack. Such attacks have targeted feature-based, example-based, and interpretable (including surrogates) XAI methods. The effectiveness of the attack is influenced by factors such as dataset type (shokriPrivacyRisksModel2021), dimension (pawelczykPrivacyRisksAlgorithmic2023; shokriPrivacyRisksModel2021), model architecture (shokriPrivacyRisksModel2021) and overfitting (pawelczykPrivacyRisksAlgorithmic2023). Some attacks have proven effective in the absence of knowledge of the training dataset or target architectures (liuPleaseTellMe2024), underscoring their practical threat potential.
@@
While interpretable models are often recommended as surrogates for explaining black-box models (mcdermidArtificialIntelligenceExplainability2021), as demonstrated by these attacks, the layer of interpretability can introduce a backdoor to the target system and lead to privacy leaks (narettoEvaluatingPrivacyExposure2022). In the example-based category, influence functions expose data instances, particularly outliers, due to their distinct characteristics and higher influence on the training process (shokriPrivacyRisksModel2021). Among feature-based methods, those using perturbations exhibit higher resilience to membership inference due to use of out-of-distribution points, however, this can also result in reduced explanation fidelity ((shokriPrivacyRisksModel2021). Conversely, feature-based methods with better explanation quality are also found to be susceptible to higher leakage (liuPleaseTellMe2024) suggesting a conflict between privacy and utility.
@@ 4.2.2 Model inversion
This category of privacy risk can result in reconstruction of datapoints partly or completely from outputs (zhangSurveyPrivacyInference2023) (Figure 7). These attacks can be conducted with black-box or white-box access to the model (fredriksonModelInversionAttacks2015; vealeAlgorithmsThatRemember2018) after it has been deployed in production systems. Attribute inference is a type of data reconstruction that can determine the values of certain attributes, generally those sensitive to individuals (yeomPrivacyRiskMachine2018) from outputs. When the model output, $f(x)$, is used for inversion, the inference model can be expressed as follows:
@@
Model inversion attacks have been documented in XAI on example-based, feature-based, and interpretable systems. shokriExploitingTransparencyMeasures2020; shokriPrivacyRisksModel2021 demonstrated a data reconstruction attack on influence functions in logistic regression models and found the attack dependent on data dimensionality. The authors designed different heuristics for low and high dimension data to improve coverage and efficiently recover more training points. goethalsPrivacyIssueCounterfactual2023 showed an explanation linkage attack using native counterfactuals generated from actual instances of the training data. The attack demonstrated the vulnerability of counterfactuals in leaking private attributes.
@@
Private images were found to be susceptible to recovery through saliency maps by zhaoExploitingExplanationsModel2021 leading to inadvertent exposure. The study found XAI systems that provided class-specific multiple explanations particularly prone to leakage. The authors also used attention transfer to highlight similar risks for non-explanation models. Other studies (dudduInferringSensitiveAttributes2022; luoFeatureInferenceAttack2022) have focused on attribute inference of tabular data using feature-based XAI. The former trained attack models using predictions and explanations to infer sensitive features. The latter used Shapley values and effectively executed the attack with limited number of queries on cloud ML services. tomaCombinationsAIModels2024 further showed that the efficacy of the proposed attack was dependent on the combination of black-box architecture and XAI method. Their findings indicate that linear models using Shapley values were particularly vulnerable to attribute inference.
@@
ferryProbabilisticDatasetReconstruction2024 designed a probabilistic white-box attack applicable to transparent models, such as decision trees and rule lists, and quantified the information about the training data contained in the model. The work found that models built using greedy algorithms leak more information compared to those built using optimal strategies. The authors also observed the attack’s capacity for misuse in launching other inference attacks such as membership and property inference.
@@
The above attacks from current literature demonstrate the potential of misuse of explanations APIs, resulting in a two-way flow (vealeAlgorithmsThatRemember2018) where inputs can be determined from outputs. Model explanations provide an effective attack surface compared to predictions (dudduInferringSensitiveAttributes2022; zhaoExploitingExplanationsModel2021), indicating the contradiction between the need for explanations in Trustworthy AI and protecting privacy (zhaoExploitingExplanationsModel2021). Data reconstruction attacks impact active users of AI systems, thus putting end-users at risk (zhaoExploitingExplanationsModel2021) and having a higher impact. In certain proposed model inversion attacks, sensitive attributes can be retrieved from models trained on non-sensitive attributes (dudduInferringSensitiveAttributes2022) while other proposed attacks demonstrate higher leakage from more important attributes (luoFeatureInferenceAttack2022) and recovery of entire training datasets (shokriPrivacyRisksModel2021). In addition, the above works highlight the misuse of XAI techniques even for models that do not provide explanations (zhaoExploitingExplanationsModel2021).
@@
A tension is also found between preserving privacy and maintaining utility of the XAI system. For instance, synthetic counterfactuals created by perturbing actual samples are shown to provide resilience to inversion in comparison to native counterfactuals. However, their usage is found to affect the plausibility and runtime of explanations (goethalsPrivacyIssueCounterfactual2023), suggesting degrading utility. Similarly, the use of multiple diverse explanations are usually recommended (voFeaturebasedLearningDiverse2023) for improving understandibility of explanations, however they are also found to contribute to leakage of privacy. Consequently restricting the access to explanation APIs has been suggested as a countermeasure (zhaoExploitingExplanationsModel2021), however such restrictions may reduce the utility to end-users.
@@ 4.2.3 Model extraction
In this attack, the target model is replicated to a significant degree of accuracy and fidelity (jegorovaSurveyLeakagePrivacy2022), thus breaching the confidentiality and the intellectual property of the model owner (Figure 8). In a typical model extraction attack, the adversary has black-box access to a deployed victim model and uses an unlabeled dataset to query it, thus building an attack dataset (yanExplanationLeaksExplanationguided2023) for cloning the model. In contrast, data-free model extraction leverages generative models to synthesize the datasets, which is advantageous when the input data is difficult to obtain (miuraMEGEXDataFreeModel2024). When the model output, $f(x)$, is utilized for this attack, the extraction of the target model function, $f^{\prime}$, can be expressed as follows:
@@
Since the extracted models can further leak personal data through membership inference and model inversion (songMachineLearningModels2017), this attack is usually used as a starting point for initiating other types of attacks (aivodjiModelExtractionCounterfactual2020; miuraMEGEXDataFreeModel2024).
@@
Model extraction attacks using explanations have been proposed across feature-based and example-based XAI methods. In the seminal work on the topic, gradient explanations, in the form of saliency maps, were found to be vulnerable by milliModelReconstructionModel2019. The use of explanations improved the attack by reducing the number of queries compared to attacks relying solely on model predictions. miuraMEGEXDataFreeModel2024 also leveraged gradient-based explanations but used data-free approach to train generative models for creating the attack dataset. The inclusion of explanations improved the quality of the generative model, and the accuracy of the cloned model improved with the diversity of the generated samples. Similarly, yanExplanationbasedDatafreeModel2023 employed data-free extraction wherein explanation loss was used to guide the generative model. Accuracy of the cloned model was improved by matching the victim model’s predictions and explanations.
@@
A different approach of extraction on feature-based XAI, used multitask learning to learn both classification and explanation tasks of the victim model (yanExplainableModelExtraction2022). Further, a model agnostic technique on gradient and pertubation-based XAI (yanExplanationLeaksExplanationguided2023), showed that explanations provided auxiliary information that enabled more efficient attacks, reducing the query budget compared to prediction-only strategies. The attack could also be applied to non-explanation models and achieved accuracy equivalent to those of explanation models.
@@
Besides the above extraction attacks targeting various feature-based XAI, from the example-based category, counterfactuals have been mainly targeted for this attack. In an extraction attack proposed by aivodjiModelExtractionCounterfactual2020, they were used to approximate the decision boundary of the victim model with high accuracy and fidelity under low query budgets. Multiple and diverse counterfactuals were found to aid the extraction process by divulging additional information to adversaries. An improvisation of the attack, to reduce the number of queries further, mitigate decision boundary shift and achieve higher agreement with the victim model, was proposed by wangDualCFEfficientModel2022. The method used the original counterfactual explanation with its own counterfactual as training pairs, to extract additional datapoints to train the cloned model. In another approach, kuppaAdversarialXAIMethods2021 proposed iterative querying of the victim model to capture the training data distribution. The method utilized distillation loss to transfer knowledge from the victim to the cloned model and was found to be successful due to the optimization of various properties such as diversity, proximity, feasibility, and sparsity.
@@
As demonstrated by the aforementioned attacks, explanation-based extraction attacks offer substantial advantages over traditional prediction-only approaches by facilitating model replication with reduced number of queries (milliModelReconstructionModel2019; miuraMEGEXDataFreeModel2024). The reduction in the number of queries benefits the adversary, especially in pay-by-query models. Certain attacks are also possible with partial knowledge of the data distribution (aivodjiModelExtractionCounterfactual2020) or in absence of overlap between the attack and training datasets (yanExplainableModelExtraction2022). In addition, in scenarios where attackers do not possess the input datasets, data-free extraction attacks are possible and the use of explanations is shown to improve the attack accuracy (miuraMEGEXDataFreeModel2024). Moreover, the diversity of the generated input datasets in such attacks is found to improve the accuracy of the cloned models (miuraMEGEXDataFreeModel2024).
@@
In addition to the direct threat to explanation models, XAI techniques can be misused for extraction of non-explanation models (yanExplanationLeaksExplanationguided2023). The use of diverse explanations, intended to build user trust in explanations, can lead to further leakage of privacy (aivodjiModelExtractionCounterfactual2020). Similarly, the optimization of counterfactuals to satisfy various properties to improve explanation quality, can reveal information to adversaries about class-specific decision boundaries thus aiding the attack (kuppaAdversarialXAIMethods2021) and leading to the conflict of explainability and privacy with utility.